### v4

# Effects of Collaboration on the Performance of Interactive Theme Discovery Systems

Abstract
NLP-assisted solutions have gained considerable traction to support qualitative data analysis. However, no unified evaluation framework exists which can account for the many different settings in which qualitative researchers may employ them. In this paper, we propose an evaluation framework to study the way collaboration settings may produce different outcomes across a variety of interactive systems. Specifically, we study the impact of synchronous vs. asynchronous collaboration using three different NLP-assisted qualitative research tools and present a comprehensive analysis of significant differences in the consistency, cohesiveness, and correctness of their outputs.
11footnotetext: Equal contribution, listed in alphabetical order.
1 Introduction
Making sense of large amounts of textual data is a common task in many academic disciplines and is generally achieved through qualitative methods such as Thematic Analysis (braun-2006) and Grounded Theory (glaser1968discovery). In these methods, researchers manually search through documents to identify common abstract themes that emerge from the data in a process commonly referred to as "inductive coding". However, manual inductive coding becomes infeasible as datasets increase exponentially in size, leading researchers to adopt Natural Language Processing (NLP) solutions to automate parts of the coding process doi:10.1146/annurev-polisci-090216-023229; IJoC10675.
In the last decade, several NLP solutions have been proposed to support qualitative research methods. These systems offer different ways to systematically wrangle large amounts of textual data to identify common patterns or themes, such as discovering hidden semantic structures using topic models smith-2018-loop; fang-2023-user, grouping documents based on similarity and broadcasting small sets of human annotations to the entire dataset pacheco-2023-interactive; chew2023llmassistedcontentanalysisusing, or providing live recommendations to users as they code the data dai-etal-2023-llm; gao-2024-collabcoder. To ensure the agency of the researchers, each system seeks to strike a balance between automation and manual coding by incorporating some type of human-in-the-loop (HitL) strategy.
Figure 1: In this study, we measure the quality of coded themes using different interactive systems under different coding configurations.
Generally, tools are evaluated to highlight the strengths and weaknesses of the proposed approach. Examples include evaluating topic coherence with and without human input fang-2023-user and contrasting machine-assisted code-book generation with a manual code-book generation dai-etal-2023-llm. In this paper, we argue that these evaluation strategies are often too narrow and do not account for the different real-world settings in which qualitative researchers employ HitL systems. In particular, we seek to answer the following questions: (1) Does collaboration setting measurably affect the quality of resulting code-books? (2) Do these findings hold across different NLP approaches?
Collaborative Qualitative Analysis (CQA) is a process where researchers work in teams to analyze a dataset to produce a consensus interpretation. Doing so can produce more reliable results, a richer understanding of the data, and insight that combines multiple perspectives cite-key. While there are several frameworks for conducting CQA, in this work we look at two common but contrasting collaboration settings. The first setting is defined by individual coders working separately before jointly consolidating their findings (asynchronous coding). The second setting is defined by coders working in groups and identifying themes through synchronous discussion (synchronous coding). Each setting has different strengths: asynchronous coding is more flexible as teams can work across different regions and time zones, whereas synchronous coding combines diverse perspectives and streamlines the annotation processes.
We test these settings on three different NLP tools for interactive inductive coding by performing experiments in the synchronous and asynchronous contexts. We choose tools that use different underlying NLP methods for this work: (1) the fang-2023-user system, which uses a human-in-the-loop topic modeling solution, (2) the pacheco-2023-interactive system, which models emerging themes with respect to their relationship to other high-level concepts which can be defined by users, and (3) the chew2023llmassistedcontentanalysisusing system, which leverages large language models (LLMs) to judge the adequacy of code definitions, as well as to broadcast them to a large dataset. By way of testing such varied approaches, we achieve a more representative understanding of how different tools perform in different collaboration settings.
To measure outcomes in synchronous vs. asynchronous collaboration, we propose a comprehensive evaluation strategy that looks at different output qualities. Namely, we are interested in measuring the consistency of outputs between synchronous and asynchronous coding, as well as the cohesiveness, and correctness of the resulting themes within each collaboration setting. Finally, we perform a small-scale user study to draw insights from the experience of the coders under each collaboration setting and make design recommendations according to our findings.
In summary, we make two main contributions: (1) We demonstrate how the collaboration setting in which qualitative research takes place affects the outcome of solutions proposed in the NLP literature, by testing three NLP-assisted inductive coding tools in synchronous and asynchronous coding contexts. (2) We propose a comprehensive evaluation strategy that measures different aspects of the outcome of NLP-assisted inductive coding tools, offering insights into the types and quality of themes that they are able to capture. We hope that our findings will inform the development of language technologies that are in line with the qualitative researcher workflows and offer insights for producing useful qualitative research tool evaluations.
2 Related Work
The overarching goal of the systems we investigate is to partially automate the qualitative coding process either by inducing topics in an interactive, semi-supervised manner fang-2023-user; smith-2018-loop, by learning user-defined themes interactively pacheco-2023-interactive; gao-2023-coaicoder, or by prompting LLMs with natural language definitions of the observed themes chew2023llmassistedcontentanalysisusing; dai-etal-2023-llm. A separate but related line of work exemplified by gao-2024-collabcoder uses LLMs to generate label recommendations as users perform the coding process. While this system is explicitly designed for asynchronous collaboration, the systems we study differ in their ability to annotate large portions of the dataset without extensive supervision.
Our research addresses a real-world use case for qualitative researchers using HiTL systems and is informed by the Human-Computer Interaction (HCI) literature jiang-2021-serendipity; feuston2021tools; chen-2018-ambiguity. A common finding in the HCI studies on computer-assisted qualitative research is the researchers’ focus on ambiguity. In traditional settings, researchers can independently code a significant portion of the dataset and meaningfully identify areas of ambiguity. However, when working with large datasets and relying on NLP systems, researchers cannot easily uncover all such areas. Solutions have been proposed to either visualize codes (drouhard-aeonium) or rank document disagreement (zade-2018-disagree) regardless of dataset size. Our evaluation differs from these works since we look at different output qualities of the resulting themes and their assignments, combining signal from group overlaps, relationships in the semantic embedding space, as well as manual post-hoc evaluations. This method of evaluation highlights areas where the coders diverge both with each other and with the model, providing another perspective on the ambiguity question.
The motivating factor for our study is the popularization of HitL systems for qualitative coding in the NLP field. With the growing number of new interactive systems comes a need for standardized evaluation across different real-world environments. Previous evaluation methods introduced with new systems have generally been ad hoc, with experiments conducted variously in group settings  choo-utopian; hoque-convisit; smith-2018-loop, on individual participants  rietz-cody, and through platforms such as MTurk  zade-2018-disagree. Our contribution provides a standardized framework for performing experiments in synchronous and asynchronous settings, as well as a set of metrics for evaluating consistency, cohesiveness, and correctness in experimental results.
3 Interactive Systems
We identify three categories of NLP techniques used in interactive systems for qualitative coding with large datasets: topic models, relational approaches, and LLMs. These techniques may be applied in a variety of ways in interactive systems, but we put special focus on their ability to help code large datasets. To maximize coverage across systems, we select a representative system from each category to use in our experiments. In this section, we briefly describe the unique aspects of each category and introduce the selected system.
Topic Models. This category includes systems that use some variation of topic modeling to find emerging themes and facilitate document assignment. These systems benefit from the relative speed of the topic model, which allow users to quickly visualize and explore the dataset. Early exploration incorporated visualizations to help users adjust parameters (Chuang2013DocumentEW), while later works implemented refinement operations that allow users to directly edit topic words and remove documents (smith-2018-loop). However, topic modeling systems are limited by their lack of malleability and predictability. Refinement operations mostly edit topic words, which can have limited impact in the final results.
We select the HitL query-driven topic model (QDTM) system introduced by fang-2023-user. The topic model is initialized by providing input queries (i.e., words that represent concepts of interest for the user) which the model uses to generate the initial topics. In our experiments, users begin by iterating through each topic and naming them based on identified themes. Users then use a set of refinement operations to edit the topic model. They can merge and split topics based on topic words, add, remove, or reorder topic words, and remove documents from topics. The next iteration of the model is only produced when the users choose to apply refinements and the prior model is saved, allowing users to return to prior iterations to test different operations. Once satisfied with the state of the topic model, the user downloads the document distribution for that iteration. Details of our experimental settings can be found in App. A.
Relational Approaches. Relational approaches combine vector semantics and structured inference to model relationships between high-level concepts. Instead of treating themes as distributions over words (as topic models do), these frameworks define themes as distributions over generalized concepts. This reflects the inductive coding process, where researchers identify patterns and concepts that are then synthesized into more abstract themes. However, their computational complexity grows with the number of dependencies considered, which hinders their ability to quickly adapt during coding sessions. Further, they rely on users to define informative concepts, making them less suited for inexperienced researchers.
We select the system introduced by pacheco-2023-interactive, which uses a two-stage relational framework. In the first stage, the system automatically partitions the dataset based on semantic similarity. The users explore each partition to identify themes, assign "good" and "bad" example documents for each theme, and input or correct supporting concepts for each example. In the second stage, the system uses the provided examples and concept relations to map the remaining dataset, only leaving documents unmapped if no theme is a sufficiently good match. The assignment procedure follows as a structured inference approach, where dependencies between concepts and themes are explicitly modeled. The unmapped documents are repartitioned as in the first stage and users are prompted to review unmapped partitions again. The process is iterated until all documents are mapped. More details about the framework and our experimental configuration can be found in App. B
Large Language Models. LLMs are ideal candidates for interactive systems, especially for tasks such as qualitative coding where the model can be prompted to produce themes or explanations without ad-hoc training (llms0shotreasoners). They have been used for theme recommendation (gao-2023-coaicoder), for code conflict resolution (gao-2024-collabcoder), and for automated document assignment (deductivecodingwithLLM). However, the flexibility of LLM outputs also leads to hallucinations, which are only partially addressed by prompt engineering. Models further suffer from biases in training which are difficult to identify and impact their ability to produce quality labels or recommendations (chen-2018-ambiguity). Additionally, their massive size is prohibitive when working with large datasets due to the high cost of inference.
We select the framework introduced by chew2023llmassistedcontentanalysisusing. In their protocol, the human coder first manually codes a representative subset of the data and drafts definitions for each code. The LLM is then prompted to label the data sample with the provided definitions. Agreement is calculated between human and model annotations using GWET’s $AC_{1}$ (gwet2008computing). The prompt is then tweaked iteratively to achieve a satisfactory level of agreement, and the best-performing version is used to prompt the model to code the rest of the dataset. Details about our experimental settings can be found in App. D.
4 Study Design
To study the effects of different collaboration settings on the performance of the three selected systems, we design a protocol that can be used for both synchronous and asynchronous settings. For each system, we conduct three asynchronous experiments with one coder each and two synchronous experiments with three coders each for a total of 15 experiments. Evaluation metrics are calculated by comparing the resulting code-books within each experimental setting (e.g. the two code-books independently created by the two synchronous groups using the topic model). The rest of this section lays out the dataset, participant demographics and experimental protocol.
Dataset. We use a common dataset for all experiments. It was originally released by pacheco-etal-2022-holistic, and it consists of approximately $85,000$ Tweets about COVID-19 vaccines. All tweets in this corpus were posted by users located in the United States, are uniformly distributed between Jan. and Oct. 2021, and contain labels for vaccination stance (e.g. pro-vax, anti-vax) and morality frames (e.g. fairness/cheating and their actor/targets.) roy-etal-2021-identifying.
Participants. We recruited a group of 20 researchers in NLP and Computational Social Science, 6 female and 14 male, between the ages of 20 and 45. This group included professors at different levels of seniority, postdoctoral researchers, and graduate and undergraduate students from two different universities. This group covers the range of researchers likely to use interactive coding systems. All participants were either well-versed in qualitative data analysis, or were explicitly trained by senior researchers to perform the task.
Coding Protocol. At the start of each experiment, participants were provided with a demonstration of all the operations in their respective systems. Every system starts with an initial partition of the data, so participants were instructed to read the first 25 samples in each partition, and manually create/name any themes they identified before freely exploring the rest of the dataset and start performing operations to find more themes.
In the topic model experiments, we suggested that participants merge and split topics based on their identified themes before making fine-grained refinements. They were then asked to refine the topic model based on their identified themes such that every topic corresponds to a unique theme. They kept re-running the model and making refinements until they were satisfied with the results, or until they failed to effect any meaningful changes.
In the relational system experiments, participants were tasked with selecting example documents for each identified theme, as well as determining concept relations for them. Following pacheco-2023-interactive, the supporting concepts considered were vaccination stances and morality frames (e.g., the identified theme “natural immunity” has an “anti-vax” stance, and is tied to the “purity” frame). Once participants were satisfied with their themes and selections, the system automatically coded the rest of the dataset. Unmapped examples were repartitioned and returned to the participants for a second (and last) round of coding.
In the experiments for the LLM-based system, participants produced natural-language definitions for each identified theme and selected a set of good examples for them. We then prompted the LLM with different task-prompt templates to find the best prompt for each set of participant-generated definitions, which was then used to code the rest of the dataset. Details of the templates, as well as the human-model agreement for the best template can be found in Appendix D.
Topic Model Relational LLM-Based
Jaccard Centroid Group Avg. Jaccard Centroid Group Avg. Jaccard Centroid Group Avg.
Sync $\bf 0.56(0.23)$ $\bf 0.98(0.05)^{**}$ $\bf 0.52(0.10)$ $\bf 0.36(0.19)$ $\bf 0.98(0.01)^{*}$ $\bf 0.52(0.07)^{*}$ $0.14(0.08)$ $0.98(0.03)$ $0.44(0.03)$
Async $0.30(0.17)$ $0.96(0.05)^{**}$ $0.51(0.09)$ $0.30(0.22)$ $0.94(0.07)^{*}$ $0.44(0.10)^{*}$ $\bf 0.17(0.11)$ $0.98(0.02)$ $\bf 0.45(0.03)$
Table 1: Avg. Consistency between Best Theme Matches across Annotator Groups. Statistically significant using a two-sample unpaired t-test with $p<0.05$. ∗∗ Near statistically significant with $p\approx 0.05$.
5 Evaluation
We use both descriptive metrics and a user study to provide a comprehensive analysis on the differences when annotating in synchronous and asynchronous settings. Our evaluation framework is comprised of three dimensions; consistency, cohesiveness & distinctiveness, and correctness, each of which uses metrics that are well-established in the literature NIPS2008_beed1360; 10.5555/3540261.3540416; pacheco-2023-interactive.
Figure 2: Two sets of annotators use a particular HiTL system to find themes. Since the same theme can be named differently by different annotators, we find the best match. In this example, the annotator 1’s theme "VaxIsBad" has been matched with annotator 2’s theme "antivax". After aligning, we calculate the similarity between these two themes using methods like Jaccard Similarity or Centroid Distance.
Consistency. Coders always run the risk of overgeneralizing or missing key themes, resulting in unclear and unsystematic results (cornish-2014-sage-cqa). We propose consistency as one dimension for measuring systematicity. Building on the idea of inter-coder reliability, we define consistency as the measure of annotators eliciting the same themes from the same set of texts (See Fig. 2). Once the dataset has been annotated using an interactive system, there are two obstacles to measuring consistency when using semi-automated annotations. First, themes with similar names may not share the same annotated documents. Similarly, themes with different names may share the same documents. We measure consistency using the overlap of documents between a given pairs of themes. To do this, we calculate the Jaccard similarity across the documents for two themes with different coders. We then select the maximum Jaccard similarity for each (unique) theme, which we consider to be the “best aligned theme”. In Fig. 2, this is shown in the middle where the colored lines indicate the best match obtained by maximum Jaccard overlap. After alignment, we calculate the average of all matched pairs for comparison across settings.
We want to account for the possibility of themes containing a different assortment of documents that otherwise remain relatively semantically similar. To do this, we also measure the semantic similarity between themes using the S-BERT embeddings of the documents assigned to them reimers-gurevych-2019-sentence. We do this in two ways: (1) Centroid similarity, where the similarity of two themes is the similarity between their two centroids, and centroids are calculated by taking the average of all document embeddings within the theme. (2) Group avg. similarity, where the similarity between two themes is the average of pair-wise cosine similarity between all documents in the two themes. As in the Jaccard metric, we report the maximum embedding similarity for each (unique) theme, which we consider to be the “best theme match”, and calculate their averages for comparison across settings.
Table 1 shows the average maximum Jaccard and embedding similarities for each experiment. We find that synchronous groups produce themes with higher average consistency for the topic modeling and relational systems. For the LLM-based system, we did not observe statistical significant differences across collaboration settings. Interestingly, the LLM-based system offered the least amount of intervention/refinement operations for users during the coding process. Whereas the other systems allowed for operations such as splitting and merging topics (topic modeling), and defining relations between concepts (relational approach), the LLM system only allowed users to alter the definition of the theme provided in the prompt. We hypothesize that the more interventions users can make, the more advantage they can take of the deliberation afforded by synchronous coding.
Topic Model Relational LLM-based
Intra-Theme Inter-Theme Intra-Theme Inter-Theme Intra-Theme Inter-Theme
All Sync $0.52(0.10)$ $0.40(0.04)$ $\bf 0.51(0.08)^{*}$ $0.42(0.05)^{*}$ $\bf 0.44(0.06)$ $0.40(0.04)$
Async $0.52(0.10)$ $0.40(0.04)$ $0.45(0.10)^{*}$ $\bf 0.34(0.11)^{*}$ $0.43(0.05)$ $\bf 0.39(0.04)$
Top 25% Sync $0.56(0.11)$ $0.39(0.05)$ $\bf 0.70(0.09)^{*}$ $0.52(0.07)^{*}$ $0.63(0.07)$ $0.55(0.05)$
Async $0.56(0.11)$ $0.39(0.05)$ $0.64(0.09)^{*}$ $\bf 0.46(0.13)^{*}$ $0.63(0.05)$ $\bf 0.54(0.05)$
Table 2: Group Avg. Similarity within Annotator Groups. Themes are considered to be more cohesive if intra-theme similarity is high and more distinctive if inter-theme similarity is low. Statistically significant using a two-sample unpaired t-test with $p<0.05$.
Figure 3: Once an annotator has identified themes and they have been propagated the full dataset, we calculate intra-theme similarity by measuring the avg. of the pairwise distances between each document within a theme (left). We calculate inter-theme similarity by measuring the avg. of pairwise distances between each document in a theme and documents assigned to all other themes (right)
Cohesiveness and Distinctiveness. Another dimension for determining the systematicity and clarity of annotation outcomes is by evaluating the similarities and differences between themes within the same code-book. We propose two metrics to measure this: cohesiveness and distinctiveness. A theme is said to be cohesive if its documents are similar to each other (measured by intra-theme similarity) and distinctive if it is dissimilar from documents in other themes within the same code-book (measured by inter-theme similarity). Intuitively, the purpose of grouping documents by theme is to create abstract representations of a dataset, where each theme represents a distinct facet of the data. If themes are not cohesive and distinctive, then it becomes hard to tell which theme a given document should belong to and the code-book falls apart.
Figure 3 shows how to evaluate these metrics for a single annotator (or annotator group). We calculate both the intra-theme similarity and the inter-theme similarity for all the themes in the code-book. Intra-theme similarity is calculated by taking the average of pair-wise similarity between all documents of the same theme. Inter-theme similarity for a given theme is calculated by taking the average pair-wise similarity of documents in that theme with documents in all other themes.
A confounding factor in these measures is that all systems provide broad coverage of documents such that even distantly related documents may be assigned to a theme. To more accurately represent the cohesiveness and distinctiveness of themes in each experiment, we perform the same calculations on a subset comprised of only the top 25% of documents most closely related to each theme. For the relational and LLM-based systems, this top quartile is selected using the distance from the centroid. For the interactive topic model, we use the weights assigned by the model.
Table 2 shows results for both the whole dataset as well as the subset of the documents closest to each theme. Overall, we find that the intra-theme similarities are always higher than inter-theme similarities, which means that themes are at least moderately cohesive and distinctive across the board. For the relational system experiments, we find that the difference between both types of metrics remains stable across collaboration modalities, and that similarities are higher in the synchronous setting. Our findings suggest that themes may be more cohesive but not more distinctive in the synchronous setting.
Results for the interactive topic model and LLM-based systems are nearly identical across collaboration modalities. Given that the operations to characterize themes offered by these two systems are substantially limited (i.e., adding and removing words for the topic model, and refining theme definitions for the LLM system), these results are not surprising. Again, we see that the wider the range of operations afforded to characterize a theme, the more advantage can be taken from synchronous deliberation.
Correctness. Interactive systems allow users to automate large portions of the coding process at the risk of producing inaccurate theme assignments. To estimate how correct the outputs of each system are, we conduct a post-hoc analysis by manually checking a randomly selected sample of $1,200$ document-theme pairs ($200$ per experimental setting). To ensure that our sample is representative of the overall dataset for each experimental setting, we split the data into quartiles based on document relatedness to each theme and select a uniform sample of themes and relatedness scores. As in the previous section, relatedness is calculated using the theme weight distribution for the interactive topic model and distance from the centroid for the the other two systems. To assess reliability, each assignment is evaluated by two annotators, with a third weighing in for tie breaks. We also find that human evaluators have moderate-to-high agreement when assessing system outputs (with an overall Krippendorff’s $\alpha$ of $0.632$), suggesting that we can trust these estimations.
Figure 4 shows the correctness results for each quartile sample per experiment. First, we observe that the relational system is not only the most accurate, but it shows negligible correctness differences in synchronous vs. asynchronous configurations. This an encouraging result, given that this system took the most advantage of synchronous deliberation based on our other metrics of quality. The other two systems showed marked differences, with the topic modeling approach producing more accurate assignments in asynchronous operation, and the LLM producing more accurate assignment in a synchronous paradigm. This observation leads us to believe that interactive systems that introduce less and weaker inductive biases from interaction are more sensitive to choices of configuration – and researchers should be aware of these possibilities when designing experiments. Second, we find that the correctness of different approaches tapers differently, based on the distance of the example from the theme. In other words, as the data samples resemble less like the human labeled ones, some models handle it more robustly than others. In our observation, relational approaches and LLM methods outperform topic models when it comes to assigning themes to distant examples.
Figure 4: Correctness w.r.t. distance from theme.
6 User Study and Recommendations
We carried out semi-structured interviews with the participants in the study. We focused on trying to understand attitudes towards the task and tools, particularly with respect to the synchronous and asynchronous modalities. Interview questions can be found in App. C. We summarize our main findings below.
Participants found working in teams eases the coding process. Teams of annotators working synchronously described their perceived advantage of working in teams – broader perspectives helped them contextualize diverse tweets easily. The annotators also valued the ability to collaboratively resolve disagreements through discussions. They cited it being easy to "break ties" quickly through conversation. In addition to easing the user experience, we also note that the consistency and cohesiveness of annotations is higher in synchronous settings. Based on this evidence, future work should consider designing systems that explicitly support synchronous deliberation.
Asynchronous coders were more likely to notice inadequacies of the tool they used. Asynchronous annotators directed most of their feedback towards the way the tools functioned. Since the bulk of the annotations were carried in isolation, difficulties navigating the tool were more salient to them than to their synchronous counterparts. Based on these findings, tool designers should focus on improving the asynchronous coding experience.
Annotators noted a lack of control and agency in the topic modeling system. Annotators expressed loss of control and agency when using the topic modeling system fang-2023-user over other systems. One annotator noted that the controls to modify topics were not satisfactory – "the merge process did not offer the ideal amount of control and made it difficult to keep track of the theme groups". Based on these findings, tool designers should consider abstracting users away from the intricacies of the underlying NLP method.
Annotators felt that starting topics often negatively impacted their annotations. Both synchronous and asynchronous annotators were dissatisfied with some of the ways in which tweets belonging to different or opposite themes were placed under the same topic in the topic modeling system. In contrast to the clustering approach used by the other two systems, which appeared to yield cleaner partitions. One of the annotators commented – "Many Anti-Vax and Pro-Vax standpoints use the same words/phrases in their tweets, which the fang-2023-user model groups together despite the stark difference in message between the two.". Some users felt frustrated at being unable to achieve the purity of topic that they hoped for, often stopping when they found that none of their changes were affecting the topics as they intended. However, one annotator did note that the benefit of using a topic modeling system was that there was some amount of theme induction done already - as opposed to the other two systems, where clusters did not have any representative words attached to them. Based on these findings, researchers should investigate ways to effect more change during interactions for topic modeling systems, as well as consider generating some initial explanations to characterize clusters for alternative approaches.
Using LLMs to conduct thematic coding at scale is expensive and less reliable. Despite the recent advances in the reasoning capabilities of LLMs, we observe that their performance at scale is not effective enough to be used for the purpose of propagating annotations. Additionally, running LLMs for the $5$ annotation configurations came with significant time and resource costs. This opens an opportunity for NLP researchers to make LLMs more reliable inductive reasoners, and to come up with prompting strategies that can allow LLMs to reliably classify documents in bulk, especially when working at scale.
Future work developing interactive systems should take into account the amount of control the system offers to the user. Quantitatively, we find that the topic modeling system produced results that were less correct across the board. This matches with our user study where users were unable to disentangle themes that had a high degree of lexical similarity. Despite being able to refine the topic model such that the highest weighted documents matched their defined themes, users had little control over whether or not the rest of the documents in the dataset were faithfully assigned.
7 Conclusion and Future Work
In this paper, we examined three categories of NLP-assisted qualitative research tools in different collaboration modalities, and conducted inductive and deductive coding on a large dataset of English texts. We designed an evaluation framework that describes the quality of the induced themes and their resulting document assignments under synchronous and asynchronous collaboration. We note that in general, the collaboration modality is a significant factor in determining if the quality of a system’s output. This is particularly true for systems like pacheco-2023-interactive, where users have a wider bandwidth of interaction that can benefit from group consensus, and provide strong inductive biases for deductive coding. We also observe that solutions based on topic modeling, although popular in the data analysis literature, can struggle with inducing cohesive themes and accurate code assignments. Finally, we show that while LLM-based solutions show promise, they pose significant challenges when it comes to coding at scale.
While this study focuses on collaboration modalities, there are numerous other variables that can affect a tool’s efficacy for qualitative coding. We believe that our proposed evaluation framework can be repurposed and expanded to evaluate a wide range of interventions, such as the underlying NLP technology, the interactive interface, the expertise and life experience of the coders, and the type of data being annotated. Through this paper, we hope to inform future work that gives way to more robust measures of an NLP tool’s capabilities as it pertains to qualitative research under realistic settings.
8 Limitations
The study presented in this paper has four main limitations.
(1) While we selected three distinct, representative tools to perform our analysis of synchronous vs. asynchronous settings, the list is of course non-exhaustive. A larger study incorporating more tools could yield additional insights.
(2) As prior work has shown 10.5555/3540261.3540416, automated metrics and human evaluation do not always correlate well for unsupervised or semi-supervised theme induction tasks like topic modeling. Although we did incorporate two types of human evaluation: a post-hoc manual verification (Sec 5) and a user study (Sec 6), both of these evaluations were limited in scope. This is primarily due to how expensive and time consuming it is to perform an exhaustive human evaluation. While we are confident in the findings of our study, a larger study could further strengthen our conclusions.
(3) We used a single dataset in our experiments, which is composed of short English texts. While we believe that this was sufficient to arrive at strong conclusions across systems and collaboration modalities, datasets with different characteristics (e.g., longer documents, non-English language) could pose additional challenges.
(4) While we look at an important variable in qualitative research settings (collaboration modality), there are several other variables that can influence the outcome of NLP-assisted solutions (e.g., choice of tool, expertise and live experience of annotators, type of data being annotated, etc.). In addition to this, we did not explore the many different consolidation strategies that are often used to bring together the perspectives of asynchronous coders. We leave the explorations of these questions for future work.
9 Ethical Considerations
To the best of our knowledge, no code of ethics was violated during the development of this project. We used publicly available tools and datasets according to their licensing agreements. For our annotation experiments, we followed IRB protocol and did not retain any personally identifiable information.
All information needed to replicate our experiments is presented in the paper. We reported all experimental settings, as well as any pre-processing steps, learning configurations, hyper-parameters, and additional technical details. Due to space constraints, some of this information was relegated to the Appendix. In addition to this, we will make the results of the annotation experiment available to the community, as well as the code to produce all of our reported results..
Appendix A Topic Model Experimental Settings
To ensure comparable results, we use the same starting distribution of 13 topics for all our experiments using the same hyperparameters as fang-etal-2021-query, which are $\alpha=1.0$, $\beta=0.5$, $\gamma=1.5$. The QDTM also allows queries to be input prior to topic model initialization to produce partitions that follow prior knowledge, but we do not take advantage of this capability. The same initial topic model is provided for all experiments.
Appendix B Experimental Settings and Interactive System Details
During the interactive coding process, researchers are provided initial clusters to identify themes. We use $K=10$ means clustering to generate initial partitions and the same partitions are provided for all experiments. Users select positive and negative examples for each theme, which are used to create a distributional representations to calculate semantic similarity with unlabeled documents. Simultaneously, supporting concepts are defined using logical rules which researchers can use to label example documents. The relationship between themes and concepts are then used as a structured inference task to predict label assignments. By using multiple sources of information, the model can much more efficiently extrapolate from a small set of manually labeled data.
B.1 Relational Approach Operational Details
Operations Description
Finding Partitions Experts can find partitions in the space of unassigned instances. We currently support the K-means Jin2010 and Hierarchical Density-Based Clustering mendelsohn-etal-2021-modeling algorithms.
Text-based Queries Experts can type any query in natural language and find instances that are close to the query in the embedding space.
Finding Similar Instances Experts have the ability to select each instance and find other examples that are close in the embedding space.
Listing Themes and Instances Experts can browse the current list of themes and their mapped instances. Instances are ranked in order of “goodness”, corresponding to the similarity in the embedding space to the theme representation. They can be listed from closest to most distant, or from most distant to closest.
Visualizing Local Explanations Experts can visualize aggregated statistics and explanations for each of the themes. To obtain these explanations, we aggregate all instances that have been identified as being associated with a theme. Explanations include wordclouds, frequent entities and their sentiments, and graphs of concept distributions.
Visualizing Global Explanations Experts can visualize aggregated statistics and explanations for the global state of the system. To do this, we aggregate all instances in the database. Explanations include theme distribution, coverage statistics, and t-sne plots JMLR:v9:vandermaaten08a.
(a) Exploratory Operations
Operations Description
Adding, Editing and Removing Themes Experts can create, edit, and remove themes. The only requirement for creating a new theme is to give it a unique name. Similarly, themes can be edited or removed at any point. If any instances are assigned to a theme being removed, they will be moved to the space of unassigned instances.
Adding and Removing Examples Experts can assign “good” and “bad” examples to existing themes. Good examples are instances that characterize the named theme. Bad examples are instances that could have similar wording to a good example, but that have different meaning. Experts can add examples in two ways: they can mark mapped instances as “good” or “bad”, or they can directly contribute example phrases.
Adding or Correcting Concepts We allow users to upload additional observed or predicted concepts for each textual instance. For instances and phrases added as “good” and “bad” examples, we allow users to add or edit the values of these concepts. The intuition behind this operation is to collect additional information for learning to map instances to themes.
(b) Intervention Operations
Table 3: Interactive Operations for the pacheco-2023-interactive System
Appendix C Semi-Structured Interview
C.1 Interviewing
We usually administered interviews after annotation sessions. In some cases, due to time constraints from the annotation session taking longer than expected, we were unable to pose these questions to the annotators. Asynchronous annotators were asked questions individually about their experience, whereas synchronous annotator groups were usually asked questions with their fellow annotators.
C.2 Script
  1. 1.
    Have you worked on annotation projects before? Did these annotation projects use qualitative coding strategies (ex: grounded theory)? How experienced are you as an annotator?
  2. 2.
    How was your experience on the COVID-19 vaccine annotation session we conducted on Sunday? Particularly, we are interested in your thoughts and feelings over the session.
  3. 3.
    You annotated in a group, working together as a team. Did you find this setup to be beneficial? What were some of the limitations you faced, both individually and and as a group, when working synchronously?
  4. 4.
    On a similar line, what would you consider to be the pros and cons if you were to annotate alone?
The last questions would be flipped based on if we are posing it to synchronous or asynchronous annotators.
Appendix D LLM-based Experimental Configuration and Prompt Details
For this study, we use the same starting partitions as used in the relational approach experiment in App. B. The original work uses existing codes from a theoretical framework whereas we use codes defined by the user during this step, but this does not affect the overall process.
code-book Gwet’s $AC_{1}$ # Unlabeled Docs
Sync 1 $0.42$ $5,548(6.5\%)$
Sync 2 $0.49$ $9,766(11.4\%)$
Async 1 $0.61$ $611(0.7\%)$
Async 2 $0.62$ $2,506(2.9\%)$
Async 3 $0.46$ $13,816(16.1\%)$
Table 4: Results for the selected prompt for each coding session using the LLM-based system. Gwet’s $AC_{1}$ is used to select the best prompt for running the full dataset. The number of unlabeled documents represent documents where the LLM produced a label not created by human annotators after running the full dataset (percentage of the dataset unlabeled).
Using the LLM, we generate 3 additional templates based on a prompt structure provided in the original work. We used the Llama 3.2 3B-Instruct (grattafiori2024llama3herdmodels) model for all generation tasks:
D.1 LLM Hyperparameters
Batch size: 32
Model: Llama 3.2 3B-Instruct
GPU: A100 40GB VRAM
Average Compute Time: 24hrs per job
Number of jobs: 5
D.2 LLM Prompts
To code this tweet, do the following:
- First, read the codebook and the tweet.
- Next, decide which code is most applicable and explain your reasoning for the coding decision.
- Finally, generate json with your code and your reason for the coding decision. The respone MUST be formatted as JSON.
Codes:
<codes>
Codebook:
<codebook>
Tweet:
<tweet>
JSON Output:
"code" : "",
"reason" : ""
To code this tweet, do the following:
First, read the codebook and the tweet.
Next, decide which code is most applicable based on the tweet’s content and explain your reasoning for the coding decision.
Finally, generate a JSON object with the selected code and provide a brief explanation for your coding decision.
The response MUST be formatted as JSON.
Codebook: Themes: <"theme": "definition">
Tweet: < "text": "<text>" >
JSON Output: < "code": "", "reason": "" >
To generate code for this tweet, provide a step-by-step explanation of how to approach the task.
First, analyze the tweet’s content and identify key concepts, such as the type of object or class being described, any specific behaviors or requirements, and relevant keywords.
Next, evaluate the codebook options and determine which one is most applicable. Explain your reasoning for your decision, including any similarities between the tweet and the code definitions, or any specific requirements mentioned in the tweet that align with a particular code.
Finally, generate a JSON object with the selected code and provide additional context, including:
A clear explanation of how you arrived at your chosen code
Any relevant notes or comments about the code’s functionality and requirements
A brief comparison to other codes in the book, if applicable
The response MUST be formatted as JSON.
Codebook: <codebook>
Tweet: <tweet>
JSON Output: < "code": "", "reasoning": "", "context": "" >
To analyze this tweet and select a relevant theme, follow these steps:
First, read the tweet and identify key concepts, such as emotions, objects, or ideas mentioned in the text.
Next, evaluate the theme options and determine which one is most applicable. Explain your reasoning for your decision, including any connections you see between the tweet’s content and the theme definitions.
Then, generate a JSON object with the selected theme and provide additional insight into your analysis. Include:
A clear explanation of how you arrived at your chosen theme
Any specific characteristics or keywords from the tweet that support your decision
A brief comparison to other themes, if applicable
The response MUST be formatted as JSON.
Themes: <Codebook>
Tweet: <tweet>
JSON Output: <"theme": "", "insight": "">